A stone that weighs 3.6 lbs has a mass of 7.9 kg. True or false
Answer #1

Solution-

As it is given stone that weighs= 3.6 lbs

so 1 lbs= 0.453592 kg

So 3.6= 3.6* 0.453592 = 1.6329312 kg

Now as we know that weight= mass*gravity

=> 1.6329312= mass* gravity

=>mass= 1.6329312/9.8

= 0.16 kg

So it is false